a. 6, 4, 1, 0, 1

b. 7, 5, 3, 3, 2, 0, 2

c. 1, -3, 6, 7, 3, 5, 5, 6, 7

d. 0, 2, 0, 0, -4, 4, -2, 4, 0, -4, 4, -4, 0, -3, -2, -4, 0, 4

I need the range, variance and standard deviation for each a, b, c and d.
Answer #1
a) Values (x) (x-mean)^2
6 12.96
4 2.56
1 1.96
0 5.76
1 1.96
Mean 2.4 SUM 25.2
Range = Highest value - lowest value
Range = 6 - 0
Range = 6.
Variance = SUM/5
Variance 25.2/5
Variance 5.04
Standard deviation = (Variance)^(.5)
Standard deviation (5.04)^(.5)
Standard deviation 2.2449944
b) Values (x) (x-mean)^2
7 14.877551
5 3.44897959
3 0.02040816
3 0.02040816
2 1.30612245
0 9.87755102
2 1.30612245
Mean 3.1428571 SUM 30.8571429
Range = Highest value - lowest value
Range = 7 - (0)
Range = 7
Variance = SUM/7
Variance 30.85714/7
Variance 4.4081633
Standard deviation = (Variance)^(.5)
Standard deviation (4.4081633)^(.5)
Standard deviation 2.0995626
c) Values (x) (x-mean)^2
1 9.67901235
-3 50.5679012
6 3.56790123
7 8.34567901
3 1.2345679
5 0.79012346
5 0.79012346
6 3.56790123
7 8.34567901
Mean 4.1111111 SUM 86.8888889
Range = Highest value - lowest value
Range = 7 - (-3)
Range = 10
Variance = SUM/9
Variance 86.88889/9
Variance 9.654321
Standard deviation = (Variance)^(.5)
Standard deviation (9.654321)^(.5)
Standard deviation 3.1071403
d) Values (x) (x-mean)^2
0 0.07716049
2 5.1882716
0 0.07716049
0 0.07716049
-4 13.8549383
4 18.2993827
-2 2.96604938
4 18.2993827
0 0.07716049
-4 13.8549383
4 18.2993827
-4 13.8549383
0 0.07716049
-3 7.41049383
-2 2.96604938
-4 13.8549383
0 0.07716049
4 18.2993827
Mean -0.277778 SUM 147.611111
Range = Highest value - lowest value
Range = 4 - (-4)
Range = 8
Variance = SUM/18
Variance 147.6111/18
Variance 8.2006173
Standard deviation = (Variance)^(.5)
Standard deviation (8.2006173)^(.5)
Standard deviation 2.863672